OpenStudy (anonymous):

Find a number that is 56 less than its square. help pls!

OpenStudy (anonymous):

Let x be the number. Then, we have: x = x^2 - 56. This is just: \[ \Large {x^2 - x - 56 = 0} \]This can easily be factored to: (x + 7)(x - 8) = 0, so x = -7 or 8 <- two numbers that satisfy the given condition. Hope that helped! :)

hero (hero):

x = x^2 - 56 x^2 - 56 - x = 0 x^2 - x - 56 = 0 x^2 -8x + 7x -7*8 = 0 x(x-8) + 7(x - 8) = 0 (x - 8)(x + 7) = 0 x - 8 = 0 x + 7 = 0 x = 8 x =-7

OpenStudy (anonymous):

@Ahaanomegas & @Hero thanks so much for your help! :D